OpenStudy (anonymous):

find f^-1(x) f(x)= ^3square root of x-5

OpenStudy (anonymous):

Do you know how to find an inverse of an equation?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Start by replacing the x and y values. and then solve for y.

OpenStudy (anonymous):

I could show you, but your equation is a little confusing.

OpenStudy (anonymous):

\[x=\sqrt[3]{y-5}\] \[x^3=y-5\] \[x^3+5=y\] \[f^{-1}(x)=x^3+5\]